# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v5

@@ 6 The Emergence of Impossibility in Transformer Architectures
To demonstrate where the impossibility may emerge in transformer architectures, let us analyze a minimal yet representative example of a tiny transformer predicting the next token in the famous English pangram:
@@
We trace the inference process, inspect every mathematical operation from input embeddings through final predictions, showing exactly where the impossibility emerges.
@@ 6.1 Model Architecture and Input Representation
$\mathbf{E}=\begin{bmatrix}0&0&0&0&0\\ 1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{bmatrix}$.
@@
It is represented by the input matix:
$\mathbf{X}_{0}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix}$,
@@
Next, we take into account the feedforward network (FFN) introducing nonlinear correction:
@@ 6.2 The Inference Process
We now analyze the inference process step by step to understand the moments the impossibility result may emerge. We shall keep things simple but realistic, recreating the essential steps of the inference process in transformer block.
@@
$\mathbf{W}_{Q}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&1\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(1)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\end{bmatrix}$, (59)
$\mathbf{W}_{Q}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1&0\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(2)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\end{bmatrix}$. (60)
@@
$\mathbf{Q}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{Q}_{3}^{(2)}=[1,0].$
Head 1 (agent) demands information stored along direction $[0,1]$. Similarly, Head 2 (agent) demands information stored along direction $[1,0]$. That means the heads develop different (orthogonal) privately known and independent attention patterns, leading to beliefs about the next token that are as different as possible.
@@
$\mathbf{K}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{K}_{3}^{(1)}=[1,0]\quad\text{and}\quad\mathbf{K}_{<3}^{(h)}=[0,0].$
@@
$\text{scores}_{3}^{(1)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707]\quad\text{and}\quad\text{scores}_{3}^{(2)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707].$
@@
$\boldsymbol{\alpha}_{3}^{(1)}=\boldsymbol{\alpha}_{3}^{(2)}=[0.25,0.25,0.50]^{\top}.$
@@
$\mathbf{V}_{\leq 3}^{(1)}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\end{bmatrix}\ \text{and}\ \mathbf{V}_{\leq 3}^{(2)}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\end{bmatrix}$.
@@
$\mathbf{V}_{3}^{(1)}=[0,3]\quad\text{and}\quad\mathbf{V}_{3}^{(2)}=[1.2,0].$
@@
$\mathbf{h}_{3}^{(1)}=[0,1.5]\quad\text{and}\quad\mathbf{h}_{3}^{(2)}=[0.6,0].$
@@
Each head then projects its context vector to residual space based on the privately held information routing knowledge encoded in $\mathbf{W}_{O}^{(h)}$. That private knowledge translates (or transmits) valuable context information into embedding vectors. Namely, each head calculates:
@@
$\mathbf{W}_{O}^{(1)}=\begin{bmatrix}0&0&0&0\\ 0&0&1&0\end{bmatrix}\quad\text{and}\quad\mathbf{W}_{O}^{(2)}=\begin{bmatrix}0&0&0&1\\ 0&0&0&0\end{bmatrix}$, (66)
@@
$\mathbf{o}_{3}^{(1)}=[0,0,0,1.5,0]\quad\text{and}\quad\mathbf{o}_{3}^{(2)}=[0,0,0,0,0.6].$
@@
$\mathbf{a}_{3}=\sum_{h=1}^{H}\mathbf{o}_{3}^{(h)}=[0,0,0,1.5,0.6]\quad\text{and}\quad\mathbf{a}_{<3}=[0,0,0,0,0],$
@@
$\mathbf{X}_{1}=\mathbf{X}_{0}+\mathbf{a}.$
@@
$\mathbf{X}_{2}=\mathbf{X}_{1}+\mathbf{W}_{2}\max(0,\mathbf{X}_{1}\mathbf{W}_{1}+\mathbf{b}_{1})+\mathbf{b}_{2}.$
That FFN operation explores context locally and introduces additional innovations in meaning, shaping the next token prediction result. It can amplify or counteract the votes injected by attention. Notice that adjustments are applied at every position separately, so that ideas provided by attention bids are interpreted and rewritten by the FFN agent.
@@ 6.2.1 Step 5: Logits and Prediction
$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6].$
@@
$\mathbf{L}_{a}=\sum_{h=1}^{H}\mathbf{l}^{(h)}=[0,0,0,0,1.5,0.6].$
That sum is combined with FFN logits $\mathbf{L}_{f}$ that introduce some creative noise. As a result, we obtain:
$\mathbf{L}=\mathbf{L}_{a}+\mathbf{L}_{f}=[0,-0.03,0.21,0.13,1.31,0.52].$
@@
$\Pi=\text{softmax}(\mathbf{L})=[0.1,0.1,0.13,0.12,\mathbf{0.38},0.17].$
@@
$\text{argmax}\ \Pi=\textit{fox}.$
@@ 6.3 Proof III: The Impossibility Theorem for Transformers
With the deep insight into the transformer inference process, we are ready to show the impossibility in that special and important setting.
@@
Since $L_{y^{*}}=\sum_{h}l^{(h)}_{y^{*}}$, we have:
$\sum_{h=0}^{H}p_{h}=\sum_{h=1}^{H}\log Z_{h}-\log Z=\Gamma.$ (70)
If heads produce non-proportional logit vectors, then $\Gamma>0$ by the strict convexity of log-sum-exp aggregation.
@@
which shows that aggregation creates excess confidence. This violates conservation, establishing impossibility. ∎
This final theorem bridges earlier impossibility results to concrete transformer architectures. It shows that concavity of scoring functions is the mathematical source of hallucination. The Jensen gap $\Gamma$ becomes a measurable quantity showing that probability aggregation violates conservation and any sufficiently capable transformer-based LLM must violate conservation of per-token information on non-trivial queries.
Let us note that the aggregate distribution cannot create Shannon information ex nihilo. Instead it reorganizes existing information to be more useful for the specific task of predicting $y^{*}$. Given fixed head logits, the Shannon entropy of a function (convex aggregate) of those logits is bounded from above by the the Shannon entropy of the fixed logits. This way information becomes more accessible, not more existent, as postulated by the reasoning with emergence operator and Theorem 5. The Jensen’s gap measures point-wise excess confidence of the predicted outcome, possibly making LLMs appear to know more (when in fact they do not).
@@ 6.4 The Hallucination Mechanics
Let us again take a closer look at the transformer micro-architecture described earlier to see how the impossibility result in Theorem 9 manifests itself in actual neural inference.
@@
$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6],$
@@
$\mathbf{L}=\mathbf{l}^{(1)}+\mathbf{l}^{(2)}+\mathbf{L}_{f}=[0,-0.03,0.21,0.13,1.31,0.52],$
@@
$\Pi=\text{softmax}(\mathbf{L})=[0.1,0.1,0.13,0.12,\mathbf{0.38},0.17].$
Therefore, the model expresses $38\%$ confidence that fox follows brown, with dog receiving only $17\%$ probability. The Jensen gap quantifies the conservation violation:
@@
where $Z_{h}=\sum_{y}\exp(l^{(h)}_{y})$ and $Z=\sum_{y}\exp(L_{y})$. We claim that this prediction constitutes hallucination in the sense defined by Theorem 9. It is a lucky one, but it is a hallucination.
To defend that claim, let us take a look at the individual head probability distributions:
$\pi^{(1)}$ $=\text{softmax}(\mathbf{l}^{(1)})=[0.105,0.105,0.105,0.105,\textbf{0.48},0.105],$ (73)
$\pi^{(2)}$ $=\text{softmax}(\mathbf{l}^{(2)})=[0.147,0.147,0.147,0.147,0.147,\mathbf{0.265}].$ (74)
Now, we can ask the following question: is there any evidence the heads can provide to justify or support the final probability
$\Pi_{fox}=0.38$ (75)
of the next token being fox?
A reasonable answer is offered by Bayesian reasoning. Each head outputs its own well-calibrated distribution $\pi^{(h)}$ conditioned by the context. Then, the aggregation process is executed that has no side information beyond what the two heads provide. Therefore, the only uncertainty to be resolved is about which head is more trustworthy for this token? But since there is no reason to prefer one head over the other, the reasonable way to resolve the uncertainty is to use a linear combination of the provided distributions a the mixture posterior:
$\hat{\Pi}=[(\pi^{(1)})^{\top},(\pi^{(1)})^{\top}]\hat{\boldsymbol{\alpha}}=\mathbf{P}\hat{\boldsymbol{\alpha}}$. (76)
One immediate solution is $\hat{\boldsymbol{\alpha}}=[1/2,1/2]^{\top}$, which does not discriminate heads. But then:
$\hat{\Pi}_{fox}=1/2\cdot(0.48)+1/2\cdot(0.147)\approx 31\%<38\%=\Pi_{fox}.$ (77)
That shows anything above $\hat{\Pi}_{fox}$ requires additional information that heads do not provide. In fact, let us see what is the best mixture we can get by calculating orthogonal projection of $\Pi$ onto the column space of $\mathbf{P}$ (see e.g. karpowicz_theory_2021 ; karpowicz2024pseudoinverseacrarc ). By taking the pseudoinverse of $\mathbf{P}$, we get the minimal norm least squares solution:
$\hat{\boldsymbol{\alpha}}=\mathbf{P}^{+}\Pi=[0.54,0.51]^{\top}\quad\text{and}\quad\hat{\Pi}_{fox}\approx 33\%.$ (78)
@@
That brings us to the conclusion that the attention mechanism violates the semantic information conservation by generating over-confident prediction of the next token. That over-confidence in response is the artifact of the softmax nonlinearity and the Jensen-gap quantified in Theorem 9.

### v7

@@ 6 The Emergence of Impossibility in Transformer Architectures
To demonstrate where and how the impossibility may emerge in transformer architectures, let us analyze a minimal yet representative example of a tiny transformer predicting the next token in the famous English pangram:
@@
We trace the inference process, inspect every mathematical operation from input embeddings through final predictions, showing exactly what is going on.
@@ 6.1 Model Architecture and Input Representation
$\mathbf{E}=\begin{bmatrix}0&0&0&0&0\\ 1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{bmatrix}$. (80)
@@
It is represented by the input matrix:
$\mathbf{X}_{0}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\end{bmatrix}$, (81)
@@
Notice that in the standard formulation of Vaswani et al. vaswani_attention_2017 , the context vectors are concatenated and then projected by a single $(Hd_{v})\times d_{m}$ matrix $\tilde{\mathbf{W}}_{O}$ to produce
$\tilde{\mathbf{a}}=[\mathbf{h}^{(1)},\mathbf{h}^{(2)},\dots,\mathbf{h}^{(H)}]\mathbf{W}_{O}.$ (84)
That is equivalent to the parallel formulation above, in which each head has an individual $d_{v}\times d_{m}$ output projection matrix $\mathbf{W}_{O}^{(h)}$,
$\mathbf{a}=\sum_{h=1}^{H}\mathbf{o}^{(h)}=\sum_{h=1}^{H}(\mathbf{h}^{(h)}\cdot\mathbf{W}_{O}^{(h)})=\tilde{\mathbf{a}}$. (85)
@@
In transformer block we also take into account the feedforward network (FFN) introducing nonlinear correction:
@@
Finally, in a more generalized (still simplified a bit) matrix form, given a query $Q$ the transformer equation then becomes:
$\begin{aligned} \boldsymbol{\alpha}^{(h)}(\mathbf{X})&=\mathrm{softmax}\left(\dfrac{1}{\sqrt{d_{k}}}\mathbf{X}\mathbf{W}^{(h)}_{Q}\mathbf{W}^{(h)\top}_{K}\mathbf{X}^{\top}\odot\mathbf{M}_{\text{casual}}\right),\\ \mathbf{X}_{k+1}&=\mathbf{X}_{k}+\sum_{h=1}^{H}\boldsymbol{\alpha}^{(h)}(\mathbf{X}_{k})\mathbf{X}_{k}\mathbf{W}^{(h)}_{V}\mathbf{W}^{(h)}_{O}\\ &+\sigma\left(\left[\mathbf{X}_{k}+\sum_{h=1}^{H}\boldsymbol{\alpha}^{(h)}(\mathbf{X}_{k})\mathbf{X}_{k}\mathbf{W}^{(h)}_{V}\mathbf{W}^{(h)}_{O}\right]\mathbf{W}^{(k)}_{1}\right)\mathbf{W}^{(k)}_{2}\\[8.61108pt] &=\mathbf{X}_{k}+\mathrm{MHA}(\mathbf{X}_{k})+\mathrm{MLP}(\mathbf{X}_{k})\quad\text{for}\quad\mathbf{X}_{0}=\mathbf{E}(Q),\\[8.61108pt] \boldsymbol{\Pi}_{k}&=\mathrm{softmax}(\mathbf{X}_{k}\mathbf{U}).\end{aligned}$ (91)
That formulation shows clearly that the transformer inference is a discrete-time dynamical system.
@@ 6.2 The Inference Process
We shall keep things simple but realistic, recreating the essential steps of the inference process in transformer block.
@@
$\mathbf{W}_{Q}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&1\\ 0&0\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(1)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\\ 0&0\end{bmatrix}$, (92)
$\mathbf{W}_{Q}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1&0\\ 0&0\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(2)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\\ 0&0\end{bmatrix}$. (93)
@@
$\mathbf{Q}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{Q}_{3}^{(2)}=[1,0].$ (94)
Head 1 demands information stored along direction $[0,1]$. Similarly, Head 2 demands information stored along direction $[1,0]$. That means the heads develop different (orthogonal) privately known and independent attention patterns, leading to beliefs about the next token that are as different as possible.
@@
$\mathbf{K}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{K}_{3}^{(1)}=[1,0]\quad\text{and}\quad\mathbf{K}_{<3}^{(h)}=[0,0].$ (95)
@@
$\text{scores}_{3}^{(1)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707]\quad\text{and}\quad\text{scores}_{3}^{(2)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707].$ (97)
@@
$\boldsymbol{\alpha}_{3}^{(1)}=\boldsymbol{\alpha}_{3}^{(2)}=[0.25,0.25,0.50]^{\top}.$ (99)
@@
$\mathbf{V}_{\leq 3}^{(1)}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\end{bmatrix}\ \text{and}\ \mathbf{V}_{\leq 3}^{(2)}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\end{bmatrix}$. (102)
@@
$\mathbf{V}_{3}^{(1)}=[0,3]\quad\text{and}\quad\mathbf{V}_{3}^{(2)}=[1.2,0].$ (103)
@@
$\mathbf{h}_{3}^{(1)}=[0,1.5]\quad\text{and}\quad\mathbf{h}_{3}^{(2)}=[0.6,0].$ (104)
@@
Each head then projects its context vector to residual space based on the privately held information. That information allows for routing knowledge encoded in $\mathbf{W}_{O}^{(h)}$ and translates (or transmits) valuable context into embedding vectors. Namely, each head calculates:
@@
$\mathbf{W}_{O}^{(1)}=\begin{bmatrix}0&0&0&0&0\\ 0&0&0&1&0\end{bmatrix}\quad\text{and}\quad\mathbf{W}_{O}^{(2)}=\begin{bmatrix}0&0&0&0&1\\ 0&0&0&0&0\end{bmatrix}$, (106)
@@
$\mathbf{o}_{3}^{(1)}=[0,0,0,1.5,0]\quad\text{and}\quad\mathbf{o}_{3}^{(2)}=[0,0,0,0,0.6].$ (107)
@@
$\mathbf{a}_{3}=\sum_{h=1}^{H}\mathbf{o}_{3}^{(h)}=[0,0,0,1.5,0.6]\quad\text{and}\quad\mathbf{a}_{<3}=[0,0,0,0,0],$ (108)
@@
$\mathbf{X}_{1}=\mathbf{X}_{0}+\mathbf{A}.$ (109)
@@
$\mathbf{X}_{2}=\mathbf{X}_{1}+\mathbf{W}_{2}\max(0,\mathbf{X}_{1}\mathbf{W}_{1}+\mathbf{b}_{1})+\mathbf{b}_{2}.$ (110)
That FFN operation explores context locally and introduces additional innovations in meaning, shaping the next token prediction result. It can amplify or counteract the votes injected by attention. Notice that adjustments are applied at every position separately, so that ideas provided by attention bids are interpreted and rewritten by the FFN agent (see Section 8).
In this example, we select the following parameters of the FFN block:
$W_{1}=\begin{bmatrix}\mathbf{I}_{d_{m}}\ \mathbf{0}_{d_{m}\times 1}\end{bmatrix}\quad\text{and}\quad W_{2}=\begin{bmatrix}0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0.13&0&0\\ 0&0&0&-0.13&-0.05\\ -0.05&0.35&0&0&0\\ 0&0&0&0&0\end{bmatrix}\quad\text{and}\quad\mathbf{b}_{i}=0.$ (111)
@@ 6.2.1 Step 5: Logits and Prediction
$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6].$ (112)
@@
$\mathbf{L}_{a}=\sum_{h=1}^{H}\mathbf{l}^{(h)}=[0,0,0,0,1.5,0.6].$ (113)
That sum is combined with input logit $\mathbf{L}_{0}$, FFN logits $\mathbf{L}_{f}$ that introduce some creative noise. As a result, we obtain:
$\mathbf{L}=\mathbf{L}_{0}+\mathbf{L}_{a}+\mathbf{L}_{f}=[0,-0.03,0.21,1.13,1.31,0.52].$ (114)
@@
$\Pi=\text{softmax}(\mathbf{L})\approx[0.09,0.08,0.11,0.26,\textbf{0.32},0.14].$ (115)
@@
$\text{argmax}\ \Pi=\textit{fox}.$ (116)
@@ 6.3 Proof III: The Impossibility Theorem for Transformers
With the deep understanding of the transformer inference process, we are ready to show the impossibility in that special and important setting.
@@
Since $L_{y^{*}}=\sum_{h}l^{(h)}_{y^{*}}$, by the strict subadditivity of the log-sum-exp function (LSE) under logit summation we have:
$\sum_{h=0}^{H}p_{h}=\sum_{h=1}^{H}\log Z_{h}-\log Z=\Gamma>0.$ (120)
@@
which shows that aggregation violates conservation, establishing impossibility. ∎
Theorem 9 bridges earlier impossibility results to concrete transformer architectures. That complementary result shows mathematical sources of hallucination in standard LLMs. The Jensen gap $\Gamma$ becomes here a measurable quantity showing that probability aggregation violates conservation and any sufficiently capable transformer‑based LLM must violate conservation of per‑token information on non‑trivial queries.
Let us note that the aggregate distribution cannot create information ex nihilo. Instead it reorganizes existing information to be more useful for the specific task of predicting $y^{*}$. This way information becomes more accessible, not more existent. The Jensen’s gap measures point-wise excess confidence of the predicted outcome, possibly making LLMs appear to know more, even when in fact they do not.
@@ 6.4 The Hallucination Mechanics
Let us again take a closer look at the example of transformer micro-architecture to see how the impossibility result in Theorem 9 manifests itself.
@@
$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6],$ (122)
@@
$\mathbf{L}=\mathbf{l}^{(1)}+\mathbf{l}^{(2)}+\mathbf{L}_{f}=[0,-0.03,0.21,0.13,1.31,0.52],$ (123)
@@
$\Pi=\text{softmax}(\mathbf{L})\approx[0.09,0.08,0.11,0.26,\textbf{0.32},0.14].$ (124)
Therefore, the model expresses $32\%$ confidence that fox follows brown, with dog receiving only $14\%$ probability. The Jensen gap quantifies the conservation violation:
@@
where $Z_{h}=\sum_{y}\exp(l^{(h)}_{y})$ and $Z=\sum_{y}\exp(L_{y})$.
We claim that prediction constitutes hallucination as defined in Theorem 9. It is a lucky one, but it is a hallucination.
To defend that claim, take a look at the individual head probability distributions:
$\pi^{(1)}$ $=\text{softmax}(\mathbf{l}^{(1)})=[0.105,0.105,0.105,0.105,\textbf{0.475},0.105],$ (126)
$\pi^{(2)}$ $=\text{softmax}(\mathbf{l}^{(2)})=[0.146,0.146,0.146,0.146,0.146,\mathbf{0.268}].$ (127)
Is there any evidence the heads can provide to justify or support the final probability of the next token being fox equal to $\Pi_{fox}=31.9\%\approx 32\%$?
A reasonable answer is offered by Bayesian reasoning. Each head outputs its own well‑calibrated distribution $\pi^{(h)}$ conditioned by the context. Then, the aggregation process is executed that has no side information beyond what the two heads provide. Therefore, the only uncertainty to be resolved is about which head is more trustworthy for this token. Since there is no reason to prefer one head over the other, one reasonable way to resolve the uncertainty is to use a linear combination of the provided distributions a the mixture posterior:
$\hat{\Pi}=[(\pi^{(1)})^{\top},(\pi^{(1)})^{\top}]\hat{\boldsymbol{\beta}}=\mathbf{P}\hat{\boldsymbol{\beta}}$. (128)
An immediate solution is $\hat{\boldsymbol{\beta}}=[1/2,1/2]^{\top}$, which does not discriminate heads. Then:
$\hat{\Pi}_{fox}=1/2\cdot(0.475)+1/2\cdot(0.146)=0.3105<0.319=\Pi_{fox}.$ (129)
Anything above $\hat{\Pi}_{fox}\approx 31\%$ requires additional information, and that is what heads do not provide.
In fact, we can see what is the best mixture we can get by calculating orthogonal projection of $\Pi$ onto the column space of individual head distributions in $\mathbf{P}$ (see e.g. karpowicz_theory_2021 ; karpowicz2024pseudoinverseacrarc ). By taking the pseudoinverse of $\mathbf{P}$, we get the minimal norm least squares solution:
$\hat{\boldsymbol{\beta}}=\mathbf{P}^{+}\Pi=[0.55,0.42]^{\top}\quad\text{and}\quad\hat{\Pi}=\mathbf{P}\hat{\boldsymbol{\beta}}=\mathbf{P}\mathbf{P}^{+}\Pi\quad\text{and}\quad\hat{\Pi}_{fox}=0.325>\Pi_{fox}.$ (130)
@@
Another insightful explanation of the hallucination mechanics in the transformer-based inference comes from the properties of logit aggregation by summation.
Theorem 10 (Attention Aggregation is Product-of-Experts).
Let $\mathbf{l}^{(h)}$ be head logits and $\pi^{(h)}=\mathrm{softmax}(\mathbf{l}^{(h)})$. With additive logit aggregation, the attention-only distribution $\Pi_{a}$ is a Product-of-Experts (PoE)—the normalized product of the head distributions:
$(\Pi_{a})_{i}=\frac{\prod_{h=1}^{H}\pi^{(h)}_{i}}{\sum_{j=1}^{|\mathcal{V}|}\prod_{h=1}^{H}\pi^{(h)}_{j}}$. (132)
Proof.
By definition of softmax, for each head $h$:
$\pi^{(h)}_{i}=\frac{\exp(l^{(h)}_{i})}{\sum_{k=1}^{|\mathcal{V}|}\exp(l^{(h)}_{k})}=\frac{\exp(l^{(h)}_{i})}{Z_{h}}$
where $Z_{h}=\sum_{k=1}^{|\mathcal{V}|}\exp(l^{(h)}_{k})$ is the partition function for head $h$. Rearranging, we have:
$\exp(l^{(h)}_{i})=Z_{h}\cdot\pi^{(h)}_{i}$
The aggregate logits corresponding to $i$-th token are given by $(\mathbf{L}_{a})_{i}=\sum_{h=1}^{H}l^{(h)}_{i}$. Therefore:
$\exp\left((\mathbf{L}_{a})_{i}\right)$ $=\exp\left(\sum_{h=1}^{H}l^{(h)}_{i}\right)=\prod_{h=1}^{H}\exp(l^{(h)}_{i})$
$=\prod_{h=1}^{H}\left(Z_{h}\cdot\pi^{(h)}_{i}\right)=\left(\prod_{h=1}^{H}Z_{h}\right)\cdot\left(\prod_{h=1}^{H}\pi^{(h)}_{i}\right)$.
As a result, the attention-only distribution is:
$(\Pi_{a})_{i}$ $=\frac{\exp\left((\mathbf{L}_{a})_{i}\right)}{\sum_{j=1}^{|\mathcal{V}|}\exp\left((\mathbf{L}_{a})_{j}\right)}$ (133)
$=\frac{\left(\prod_{h=1}^{H}Z_{h}\right)\cdot\left(\prod_{h=1}^{H}\pi^{(h)}_{i}\right)}{\left(\prod_{h=1}^{H}Z_{h}\right)\cdot\sum_{j=1}^{|\mathcal{V}|}\left(\prod_{h=1}^{H}\pi^{(h)}_{j}\right)}$ (134)
$=\frac{\prod_{h=1}^{H}\pi^{(h)}_{i}}{\sum_{j=1}^{|\mathcal{V}|}\prod_{h=1}^{H}\pi^{(h)}_{j}}$, (135)
which is the Product-of-Experts (PoE) formula. ∎
In other words, the final probability of each token is proportional to the product of its probabilities across all heads. The multiplicative structure, induced by the logit aggregation, strongly rewards consensus and severely penalizes disagreement. In particular, if any head assigns near-zero probability to a token, the aggregate probability of PoE will also be near-zero (veto-power).
That makes the aggressive Product-of-Experts (PoE) fundamentally different from the smoothing Mixture-of-Experts (MoE) aggregation of bids. The PoE introduces the focusing effect, or excess confidence into the inference process. The tokens for which consensus exists across heads (the product is relatively large) receive a large boost (beyond evidence), forcing the probability to focus excessively on them, while reducing it for other tokens.
In addition, as we have seen earlier, the full transformer block introduces bias from the input residual and the FFN output in the aggregate logit vector. That bias modifies the token probabilities independently of the head bids, potentially amplifying or dampening the head PoE result. Consequently, the attention block fabricates certainty of knowledge beyond the constituent evidence.
That violates the truthfulness and semantic information conservation by generating unjustified predictions, as stated in Theorem 9.